OpenStudy (anonymous):

simplify x^1/3 times x^2/5

OpenStudy (mertsj):

Remember, if you are doing a multiplication problem and the bases are the same, then add the exponents.

jimthompson5910 (jim_thompson5910):

Add the exponents to get 1/3 + 2/5 = 5/15 + 6/15 = 11/15 So the answer is \[\Large x^{\frac{11}{15}}\]

OpenStudy (anonymous):

but ytou cant add fractions if the demoniators arent the same

OpenStudy (anonymous):

you can add any fractions regarless of the denominators.... to add them just get both fractions to their equivalent form such that the denominators are the same.